OpenStudy (anonymous):

Simplify the rational expression 3r^2-9r --------- r-3 Identify any excluded values.

OpenStudy (anonymous):

\[\LARGE {3r^2-9r\over r-3}={3r^2-3^2r\over r-3}={3r(r- 3)\over r-3}=3r\]

OpenStudy (anonymous):

Can you show how you got those numbers? I'm pretty lost.

OpenStudy (anonymous):

you just put those numbers in

OpenStudy (anonymous):

Which number's where I don't understand what your telling me :o

OpenStudy (anonymous):

ok here's what I've done: \[\LARGE 9=3^2\] right? and from \[\LARGE (3r^2-9r)=(3r^2-3^2r)\] are we clear till here? :)

OpenStudy (anonymous):

Yes! Thank you so much :D

OpenStudy (anonymous):

ok hold on... now we have: from: \[\LARGE (3r^2-3^2r)=3r(r-3)\] because they are equal, if we multiply them out, it becomes as it was in the beginning... \[\LARGE 3r(r-3)=3r\cdot r-3r\cdot 3=3r^2-3^2r \] so once we have: \[\LARGE {3r(r-3)\over r-3}=3r\cdot {r-3 \over r-3}=3r\cdot 1=3r\] do you get it ? :)

OpenStudy (anonymous):

Yes. I get it, thank you for explaining this to me.

OpenStudy (anonymous):

You're welcome :)